OpenStudy (anonymous):

Find f'(x) then find the tangent line to the graph of y=f(x) at x=a. Please see below for full formula

OpenStudy (anonymous):

Use \[\lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h }\] to find f'(x) and then find the tangent line to the graph of y=f(x) at x=a.

OpenStudy (anonymous):

i think the answer would depend largely on \(f(x)\) and also on \(a\)

OpenStudy (anonymous):

Oh, yea and \[f(x)=\sqrt{x+1}\] a=8

OpenStudy (anonymous):

write \[\lim_{h\to 0}\frac{\sqrt{8+1+h}-\sqrt{8+1}}{h}\] as a first step then \[\lim_{h\to 0}\frac{\sqrt{9+h}-3}{h}\] as a second step

OpenStudy (anonymous):

the h just goes into the square root like that? and i thought you have to keep the x in the equation. confused.

OpenStudy (anonymous):

then multiply top and bottom by the conjugate of \(\sqrt{9+h}-3\) which is \(\sqrt{9+h}+3\)

OpenStudy (anonymous):

you are told what \(x\) is, it is \(8\) so i replaced \(x\) by \(8\)

OpenStudy (anonymous):

you can do it with \(x\) if you like, and then plug in the 8 later you will get the same answer

OpenStudy (anonymous):

but first i need to find the equation of the tangent line, so can you keep the x in there. and im more confused about the h.

OpenStudy (anonymous):

ok what do you need to find an equation of a line?

OpenStudy (anonymous):

you need the formula. lol

OpenStudy (anonymous):

actually you need two pieces of information a point and a slope we can find the point easy enough, since you are told \(a=8\) and \(f(x)=\sqrt{x+1}\) making \(f(8)=\sqrt{8+1}=3\) and the point is therefore \((8,3)\)

OpenStudy (anonymous):

the second piece of information you need is the slope that is what we are finding now

OpenStudy (anonymous):

ok, so like will this be the starting point?